ABC Ltd shares are assumed to be trading at their fair value of $49.02. The dividend per share next year is $3.98 and grows at a constant rate each year. The appropriate discount rate is 16% p.a.. What is the expected price of ABC shares 5 year from now? (round to nearest cent)
Answer #1
1) Growth rate implied in the current price is to be
found out.
As per the constant dividend growth formula
P0 = D1(r-g), where D1 = next expected dividend,
r = required return or discount rate and g = growth
rate in dividends.
Substituting values we have,
49.02 = 3.98/(0.16-g)
0.16-g = 3.98/49.02
0.16-3.98/49.02 = g = 7.88%
2) The expected price per the constant dividend growth model at EOY 5 = D6/(r-g)
Where, D6 = Expected divid,end for the 6th year
r = required return or discount rate and g = growth
rate.
Substituting values into the above equation
Price at EOY 5 = 3.98*(1.0788^4)/(0.16-0.0788) = $          66.39